OpenStudy (anonymous):

what does it mean for a number to be the least common multiple of two numbers

OpenStudy (anonymous):

If your two numbers are x, and y; the lowest common multiple is the lowest number for which \[n*x=m*y \text{ for some }n,m\in \mathbb{N}\]

OpenStudy (anonymous):

The least common multiple is the smallest number that both of your original numbers can "reach, or multiply to. For example, between the numbers 6 and 4, you can multiply them to reach 24, which is a multiple of both numbers. However, it is the not the least, as 6 times 2 is 12, while 4 times 3 is 12. So the Least Common Multiple is 12. What I do is multiply the two numbers, then divide by half and see if that number is a multiple of your two original numbers. Keep dividing by 2 until you reach a number that is not a multiple of both original numbers, or just doesn't make sense.